OpenStudy (anonymous):

solve using elimination method 0.3x-0.2y=4 0.5x+0.5y=-15/19

OpenStudy (anonymous):

panther77 didn't we just work on this one earlier?

OpenStudy (anonymous):

i thought it was a different one but i may be wrong sorry

OpenStudy (anonymous):

no it's the same.

OpenStudy (anonymous):

i'll eliminate y in order to get the value of x you can also elimiate the x first. so i multiplied the first equation by 5 and multiplied the second equation by 2 1.5x - y = 20 x+ y = -30/19 ----------------- + 2.5x = 20 - 30/19 2.5x= (380-30)/19 2.5x = 350/19 x = 140/19 then plug the value of x into one of those equations to get the value of y

OpenStudy (anonymous):

im sorry how would i figure y im unsure with the fractions?

OpenStudy (anonymous):

put the value of y to this equation 0.3x-0.2y=4 you can substitute x with 140/19, as we've got x from elimination before. 0.3(140/19) - 0.2y=4 or this 0.5x+0.5y=-15/19 0.5(140/19) - 0.5y = -15/19 yeah the fraction makes the calculation difficult. are you sure the second equation is 0.5x+0.5y=-15/19? why dont you check your question again...

OpenStudy (anonymous):

yes that is correct

OpenStudy (anonymous):

ok, then it's just a matter of calculation 0.5(140/19) - 0.5y = -15/19 -0.5y = (-15/19) - (70/19) -0.5y=-85/19 y=170/19